Write the equation of the line that passes through point (5, 19) with a slope of 5.
Answer by checkley75(3666)

Y=mX+b (m)=SLOPE, (b)=Y INTERCEPT
19=5*5+b
b=19-25
b=-6 THUS THE EQUATION IS
Y=5X-6